# Age-Optimal Sampling and Routing under Intermittent Links and Energy Constraints

### v1

@@ IV-B Bounds on the Optimal Average AoI
$\lambda^{\star}$ is upper and lower bounded by:
$0\leq\lambda^{\star}\leq\min_{k\in\mathcal{R}_{\infty}}\frac{3\mu_{k}+w_{k}}{2}+\frac{\sigma_{k}^{2}}{2(\mu_{k}+w_{k})}=\lambda^{u},$ (36)
where
$w_{k}=\left(\frac{C_{s}+G_{k}\mu_{k}}{E_{\max}}-\mu_{k}\right)^{+}.$ (37)
@@ V Numerical Solutions
In the inner layer, for a fixed pair $(\lambda,c)$, we approximate the auxiliary function $h(\lambda,c)$ defined in (12) through the Average Cost Optimality Equation (ACOE) (30). In the outer layer, we update $(\lambda,c)$ iteratively until the Dinkelbach equilibrium $h(\lambda^{\star},c^{\star})=0$ and the energy constraint are simultaneously satisfied.
@@ V-A1 Challenge 1: Hybrid Action Space
The first challenge in computing $h(\lambda,c)$ arises from the hybrid nature of the action space, where sampling actions $z\in\mathbb{R}^{+}$ are continuous, while routing actions $r\in\mathcal{R}$ are discrete and depend on the current route availability. By leveraging (31), which gives the optimal sampling policy
$z^{\star}(y,\mathbf{l},r;\lambda,c)=(\lambda+cE_{\max}-\mu_{r}-y)^{+},$
we substitute $z^{\star}(\cdot)$ into the ACOE (30). This transformation yields a simplified SMDP with a countable routing action space:
@@ V-A2 Challenge 2: Uncountable State Space
The second challenge stems from the uncountable nature of the continuous state variable $y\in\mathbb{R}^{+}$. Evaluating (38) over the entire state space is computationally intractable. Traditional discretization methods approximate $\mathcal{S}$ by a finite grid of $M$ points $\{y_{1},\ldots,y_{M}\}$, introducing a quantization error $\epsilon_{M}$ that vanishes asymptotically but increases computational cost quadratically with $M$. The overall complexity of the Relative Value Iteration (RVI) approach scales as $\mathcal{O}(NM^{2})$, where $N=|\mathcal{N}|$.
@@ V-B Proposed ReaVI Algorithm with Energy Constraint
Algorithmic Solution: The resulting ReaVI algorithm iteratively updates $\{G(r;\lambda,c)\}_{r\in\mathcal{N}}$ and $h(\lambda,c)$ until convergence. In the middle layer, the Dinkelbach update for $\lambda$ is done via bisection search for a fixed $c$. In the outer layer, a bisection search for $c^{\star}$ is done to enforce the energy constraint. Given $(\lambda,c)$ and a solution $\{G(r;\lambda,c)\}$, the selection of the one-step route is:
@@
Input: $\lambda,\{\tau_{a,b}\}_{a\neq b,\;a,b\in\mathcal{N}},\mathbf{l},\{G(r;\lambda,c)\}_{r\in\mathcal{N}:\;l_{r}=0}$
@@ 1 Initialize:
$i\leftarrow 1$ ;
// Iteration number initialization
@@
$a_{\text{old}}(\mathbf{l})\leftarrow a_{1}(\mathbf{l})$ ;
// Store initial value
2 repeat
3 $\tau_{i}(\mathbf{l})\leftarrow\min_{r>a_{\text{old}}(\mathbf{l})}\tau_{a_{\text{old}}(\mathbf{l}),r}$;
4 $a_{i+1}(\mathbf{l})=\arg\min_{r>a_{\text{old}}(\mathbf{l})}\tau_{a_{\text{old}}(\mathbf{l}),r}$;
 $a_{\text{old}}(\mathbf{l})\leftarrow a_{i+1}(\mathbf{l})$ ;
 // Update newly found optimal route
 $i=i+1$ ;
 // Set for next iteration
@@ 4 $a_{i+1}(\mathbf{l})=\arg\min_{r>a_{\text{old}}(\mathbf{l})}\tau_{a_{\text{old}}(\mathbf{l}),r}$;
6until $a_{i}(\mathbf{l})=\max_{r}l_{r}:l_{r}=0$;
Output: $\{\tau_{j}(\mathbf{l)}\}_{j\in\{1,\ldots,i-1\}}$ and $\{a_{j}(\mathbf{l)}\}_{j\in\{1,\ldots,i\}}$
@@ 27   Run Algorithm 1 and find the policy $r^{\star}(y,\mathbf{l})$ in (44);
32
// Mixing to hit the boundary
33 with $c^{-}<c^{+}$, compute $q\leftarrow\frac{E_{\max}-E(\lambda,c^{-})}{E(\lambda,c^{+})-E(\lambda,c^{-})}\in[0,1]$;
34 Randomize between the two stationary policies for $c^{+}$ and $c^{-}$, $\pi_{c^{+}}$ and $\pi_{c^{-}}$ respectively. $\pi^{\star}$ uses $\pi_{c^{+}}$ with probability $q$ and $\pi_{c^{-}}$ with probability 1q1-q
@@ VI Proof of Optimal Threshold Structures
Given a specific route $r$ for $Q(y,\mathbf{l},z,r;\lambda,c)$, we first solve the conditionally optimal $z^{\star}(y,\mathbf{l};r,\lambda,c)$.
  • Case 1: If $\lambda+cE_{\max}-\mu_{r}-y\leq 0$, we have that
    $\frac{\partial Q(y,\mathbf{l},z,r;\lambda,c)}{\partial z}=z+y+\mu_{r}-\lambda-cE_{\max}\geq 0.$ (46)
    In this case, $Q(y,\mathbf{l},z,r;\lambda,c)$ is monotonically increasing with $z$ given a specific $r$ and $y$, which indicates that $z^{\star}(y,\mathbf{l};r,\lambda,c)=0$.
  • Case 2: If $\lambda+cE_{\max}-\mu_{r}-y>0$, from (46) we can establish that if $z\in(0,\lambda+cE_{\max}-\mu_{r}-y)$, $Q(y,\mathbf{l},z,r;\lambda,c)$ is monotonically decreasing with $z$; if $z\in[\lambda+cE_{\max}-\mu_{r}-y,\infty)$, $Q(y,\mathbf{l},z,r;\lambda,c)$ is monotonically increasing with $z$. As a result, $z^{\star}(y,\mathbf{l};r,\lambda,c)=\lambda+cE_{\max}-\mu_{r}-y$.
Combining the aforementioned two cases yields:
$z^{\star}(y,\mathbf{l};r,\lambda,c)=(\lambda+cE_{\max}-\mu_{r}-y)^{+}.$ (47)
@@
$Q(y,\mathbf{l},r)=$ $-\frac{\left((\lambda^{\star}+cE_{\max}-\mu_{r}-y)^{+}\right)^{2}}{2}$ (48)
$+(y+cG_{r}-cE_{\max}-\lambda^{\star})\mu_{r}+cC_{s}$
$+\frac{\sigma_{r}^{2}+\mu_{r}^{2}}{2}+\mathbb{E}_{Y^{\prime}\sim Q_{r},\mathbf{l}^{\prime}\sim\boldsymbol{p}}\big[V^{*}(Y^{\prime},\mathbf{l}^{\prime};\lambda^{\star})\big].$
With the notation $Q(y,\mathbf{l},r)$, the ACOE turns to:
@@
To analyze the threshold structure of $r^{\star}(y,\mathbf{l})$, the following lemmas discuss some important properties of $Q(y,\mathbf{l},r)$ and $V^{*}(y,\mathbf{l};\lambda^{\star},c)$.
@@
The action-value function $Q(y,\mathbf{l},r)$ is independent of the state $\mathbf{l}$, i.e
$Q(y,\mathbf{l},r)=Q(y,r),\quad\forall\mathcal{R}.$ (51)
Hence,
$V^{*}(y,\mathbf{l};\lambda^{\star},c)=\min_{r\in\mathcal{R}}\{Q(y,r)\}.$ (52)
@@
Even though $V^{*}(y,\mathbf{l};\lambda^{\star})$ depends on $\mathbf{l}$, once route $r$ has been selected, the term $\mathbb{E}_{Y^{\prime}\sim Q_{r},\mathbf{l}^{\prime}\sim\boldsymbol{p}}\big[V^{*}(Y^{\prime},\mathbf{l}^{\prime};\lambda^{\star},c)\big]$ in the right-hand side of (48) is independent of $\mathbf{l}$. This completes the proof. ∎
@@
The following assertions hold true:
@@
  3. 3.
    For any routes $j,k$ such that $\mu_{j}>\mu_{k}$, we have
    $\frac{\partial Q(y,j)}{\partial y}\geq\frac{\partial Q(y,k)}{\partial y},\forall y\in\mathbb{R}^{+}.$ (53)
@@
With (1) and (3) of Lemma 5 in hand, we can then establish the following lemma, which indicates that the optimal routing policy $r^{\star}(y,\mathbf{l})$ is monotonically non-decreasing with $y$:
@@
Consider $N$ routes with their mean delays satisfying $\mu_{1}\geq\mu_{2}\cdots\geq\mu_{N}$, if route $j$ is optimal at $y=y^{\star},\mathcal{R}$, we have that
$r^{\star}(y,\mathbf{l})=\operatorname*{arg\,min}_{r\in\mathcal{R}(\mathbf{l})}\{Q(y,r)\}\geq j,\text{if }y>y^{\star},$ (54)
$r^{\star}(y,\mathbf{l})=\operatorname*{arg\,min}_{r\in\mathcal{R}(\mathbf{l})}\{Q(y,r)\}\leq j,\text{if }y<y^{\star}.$
@@
See Appendix C
[…]
$z^{\star}(y)=z^{\star}(y;r^{\star}(y,\mathbf{l}),\lambda^{\star},c)=(\lambda^{\star}+cE_{\max}-\mu_{r^{\star}(y,\mathbf{l})}-y)^{+}.$ (55)
For a given constant-value interval $\big[\tau_{k-1}(\mathbf{l}),\tau_{k}(\mathbf{l})\big)$ where $r^{\star}(y,\mathbf{l})=a_{k}(\mathbf{l})$, the optimal sampling policy is defined as:
$z^{\star}(y)=({\lambda^{\star}+cE_{\max}-\mu_{a_{k}(\mathbf{l})}}-y)^{+}.$ (56)
Let $\beta_{k}^{\star}(\mathbf{l})\triangleq\lambda^{\star}-\mu_{a_{k}(\mathbf{l})}$, we next prove that the water-filling levels $\beta_{k}^{\star}(\mathbf{l})$ are strictly increasing with the index $k$. First, we can show that for any $i<j$, it follows that $\mu_{a_{i}(\mathbf{l})}\leq\mu_{a_{j}(\mathbf{l})}$, which leads to
$\beta_{i}^{\star}(\mathbf{l})$ $=(\lambda^{\star}+cE_{\max}-\mu_{a_{i}(\mathbf{l})})^{+}\leq(\lambda^{\star}+cE_{\max}-\mu_{a_{j}(\mathbf{l})})^{+}$ (57)
$=\beta_{j}^{\star}(\mathbf{l}).$
Next, we prove that $\beta_{i}^{\star}(\mathbf{l})\neq\beta_{j}^{\star}(\mathbf{l})$ for $i\neq j$. This is achieved by the following lemma, which indicates that $\mu_{a_{i}(\mathbf{l})}\neq\mu_{a_{j}(\mathbf{l})}$ for $i\neq j$.
@@
Let $\mathcal{R^{\star}}=\big\{a_{1}(\mathbf{l}),\ldots,a_{K+1}(\mathbf{l})\big\}$ denote the set of routes used by the age-optimal policy for a given $\mathbf{l}$ and let $\mathcal{G}_{\mu}$ be defined as
$\mathcal{G}_{\mu}\triangleq\{r\in\mathcal{R}(\mathbf{l}):\mu_{r}=\mu\}.$ (58)
Then, at most one route from $\mathcal{G}_{\mu}$ can belong to the optimal set $\mathcal{R}^{\star}$:
$|\mathcal{R^{\star}}\cap\mathcal{G}_{\mu}|\leq 1,\quad\forall\mu\in\mathbb{R}^{+}.$ (59)
@@
With lemma 7 and (57), we establish that $\beta_{1}(\mathbf{l})^{\star}<\cdots<\beta_{K+1}^{\star}(\mathbf{l})$.
Finally, we prove that there are at most $\frac{N(N-1)}{2}$ unique routing thresholds. For $a<b$, we define $\tau_{a,b}$ as where the equality $Q(\tau_{a,b},a)=Q(\tau_{a,b},b)$ is satisfied. By Lemma 6, $\tau_{a,b}$ exists if and only if $Q(0,b)\leq Q(0,a)$. Then, if the inequality
$Q(0,1)\leq Q(0,2)\leq\dots\leq Q(0,N)$ (60)
is satisfied, $\tau_{a,b}$ exists $\forall a,b\in\mathcal{N},\;a\neq b$. Hence, there are at most $\binom{N}{2}=\frac{N(N-1)}{2}$ unique routing thresholds. Note that the existence of $\tau_{a,b}$ does not necessitate its presence in the optimal solution.
VII Special Case: $\boldsymbol{p}=\boldsymbol{1}$, $E_{\max}=\infty$
This section focuses on the case where all routes are always available, i.e $\boldsymbol{p}=\boldsymbol{1}$, and there is no energy constraint upon the system. Since all routes are available at each interval, availability is no longer a part of the state space. At each decision instance a route is picked from the set $\mathcal{N}$. This problem reduces to the problem in [Mobihoc25]. Here, we present the differences of the optimal solution from problem 1.
Since $\mathcal{R}(\mathbf{l})=\mathcal{N}$ for $\forall\mathbf{l}$, the maximum number of routing thresholds is reduced to N1N-1. However, due to the independence of the waiting time to the availability probabilities $\boldsymbol{p}$, the maximum number of waiting thresholds is $N$.
The optimal routing action at the $i$-th epoch is given by
$R_{i}^{\star}=\sum_{k=1}^{K+1}(a_{k}-a_{k-1})u(Y_{i}-\tau_{k-1}),$ (61)
where $K\leq N-1$, $\tau_{0}\triangleq 0$, $a_{0}\triangleq 0$, and $u(t)$ is the unit step function.
The optimal waiting time at the $i$-th epoch has a modified expression where the energy constraint is removed from (29) and is given by
$Z_{i}^{\star}=\left(\lambda^{\star}-\mu_{R_{i}^{\star}}-Y_{i}\right)^{+}.$ (62)
Lemma 2 turns into
$\beta_{k}^{\star}<\tau_{k}.$ (63)
The upper bound to $\lambda^{\star}$ can be updated as $\min_{i}\left\{\frac{3\mu_{i}}{2}+\frac{\sigma_{i}^{2}}{2\mu_{i}}\right\}$ since all routes are available at all times.
VIII Simulation Results

### v2

@@ IV-B Bounds on the Optimal Average AoI Penalty
$\lambda^{\star}$ is bounded by:
$f(0)\leq\lambda^{\star}$ (25)
$\leq\min_{k\in\mathcal{R}_{\infty}}\left(\frac{\mathbb{E}_{Y,Y^{\prime}\sim Q_{k}}[F(Y^{\prime}+w_{k}+Y)-F(Y^{\prime})]}{\mu_{k}+w_{k}}\right)=\lambda^{u},$
where $F(t)\triangleq\int_{0}^{t}f(\tau)d\tau$, $Y$ and $Y^{\prime}$ are independent random variables representing consecutive delays on route $k$, and $w_{k}$ is the minimum constant waiting time required to satisfy the energy constraint on route $k$: $w_{k}=\left(\frac{C_{s}+G_{k}\mu_{k}}{E_{\max}}-\mu_{k}\right)^{+}$.
@@
Remark: The upper bound corresponds to the performance of a naive stationary policy that exclusively uses route $k$ with a constant waiting time $w_{k}$. In the special case of linear AoI ($f(t)=t$), the numerator simplifies to $\frac{\mathbb{E}[(Y^{\prime}+w_{k}+Y)^{2}]-\mathbb{E}[Y^{\prime 2}]}{2}$, recovering the explicit bound $\frac{3\mu_{k}+w_{k}}{2}+\frac{\sigma_{k}^{2}}{2(\mu_{k}+w_{k})}$.
@@ V Numerical Solutions
In the inner layer, for a fixed pair $(\lambda,c)$, we approximate the auxiliary function $h(\lambda,c)$ defined in (10) through the Average Cost Optimality Equation (ACOE) (19). In the outer layer, we update $(\lambda,c)$ iteratively until the Dinkelbach equilibrium $h(\lambda^{\star},c^{\star})=0$ and the energy constraint are both simultaneously satisfied.
@@ V-A1 Challenge 1: Hybrid Action Space
The first challenge in computing $h(\lambda,c)$ arises from the hybrid nature of the action space, where sampling actions $z\in\mathbb{R}^{+}$ are continuous, while routing actions $r\in\mathcal{R}$ are discrete and depend on the current route availability. By leveraging (20), which gives the optimal sampling policy $z^{\star}(y,\mathbf{l},r;\lambda,c)=(\beta^{\star}(r,\lambda,c)-y)^{+},$ we substitute $z^{\star}(\cdot)$ into the ACOE (19). This transformation yields a simplified SMDP with a countable routing action space:
@@ V-A2 Challenge 2: Uncountable State Space
The second challenge stems from the uncountable nature of the continuous state variable $y\in\mathbb{R}^{+}$. Evaluating (26) over the entire state space is computationally intractable. Traditional discretization methods approximate $\mathcal{S}$ by a finite grid of $M$ points $\{y_{1},\ldots,y_{M}\}$, introducing a quantization error $\epsilon_{M}$ that vanishes asymptotically with $M$. The overall complexity of the Relative Value Iteration (RVI) approach scales as $\mathcal{O}(|\mathcal{N}|M^{2})$.
@@ V-B Proposed ReaVI Algorithm with Energy Constraint
Algorithmic Solution: The resulting ReaVI algorithm iteratively updates $\{G(r;\lambda,c)\}_{r\in\mathcal{N}}$ and $h(\lambda,c)$ until convergence. In the middle layer, the Dinkelbach update for $\lambda$ is done via bisection search for a fixed $c$. In the outer layer, a bisection search for $c^{\star}$ is done to enforce the energy constraint. Given $(\lambda,c)$ and $\{G(r;\lambda,c)\}$, the selection of the one-step route is:
@@
Input: $\lambda,\mathbf{l},\{\mathcal{T}_{a,b}\}_{a\neq b},\{G(r;\lambda,c)\}_{r\in\mathcal{N}:\;l_{r}=0}$
@@ 1 Initialize:
2 $i\leftarrow 1$;
3 $\tau_{0}(\mathbf{l})\leftarrow 0$;
@@
4 $a_{\text{old}}\leftarrow a_{1}(\mathbf{l})$;
@@
6while true do
 // Find the set of all future potential switching points
7 $\mathcal{P}\leftarrow\{(t,r):r\in\mathcal{R}(\mathbf{l}),r\neq a_{\text{old}},t\in\mathcal{T}_{a_{\text{old}},r},\;t>\tau_{i-1}(\mathbf{l})\}$;
8
9 if $\mathcal{P}=\emptyset$ then
    Break ;
    // No more crossovers; $a_{\text{old}}$ is optimal for $y\in[\tau_{i-1},\infty)$
10
11
 // Identify the very next threshold and the route to switch to
12 $\tau_{i}(\mathbf{l})\leftarrow\min\{t:(t,r)\in\mathcal{P}\}$;
13 $a_{i+1}(\mathbf{l})\leftarrow\text{route }r\text{ corresponding to }\min\mathcal{P}$;
14
15 $a_{\text{old}}\leftarrow a_{i+1}(\mathbf{l})$;
16 $i\leftarrow i+1$;
17
Output: $\{\tau_{j}(\mathbf{l)}\}_{j=1}^{i-1}$ and $\{a_{j}(\mathbf{l)}\}_{j=1}^{i}$
@@ 27   Run Algorithm 1 and find the policy $r^{\star}(y,\mathbf{l})$ in (32);
32Compute the average epoch lengths $\bar{T}^{+}$ and $\bar{T}^{-}$ of $\pi_{c^{+}}$ and $\pi_{c^{-}}$;
33 Compute target time-fraction $\alpha\leftarrow\frac{E_{\max}-E(\lambda,c^{-})}{E(\lambda,c^{+})-E(\lambda,c^{-})}\in[0,1]$;
34 Compute randomization probability: $q\leftarrow\frac{\alpha\bar{T}^{-}}{\alpha\bar{T}^{-}+(1-\alpha)\bar{T}^{+}}$ Define $\pi^{\star}$ as choosing $\pi_{c^{+}}$ with probability $q$ and $\pi_{c^{-}}$ with probability 1q1-q at the start of each cycle;
35
@@ VI Proof of Optimal Threshold Structures
Given a specific route $r$, we determine the conditionally optimal waiting time $z^{\star}(y,\mathbf{l};r,\lambda,c)$ by analyzing the partial derivative of $Q$ with respect to $z$. Differentiating the objective function with respect to $z$ yields the expected marginal cost of waiting:
$\frac{\partial Q(y,\mathbf{l},z,r;\lambda,c)}{\partial z}=\mathbb{E}_{Y^{\prime}\sim Q_{r}}[f(y+z+Y^{\prime})]-(\lambda+cE_{\max}).$ (34)
Since $f(\cdot)$ is a monotonic non-decreasing function, the term $\mathbb{E}_{Y^{\prime}}[f(y+z+Y^{\prime})]$ is monotonically non-decreasing in $z$. We can therefore identify the optimal waiting time by comparing the expected penalty against the cost threshold $\lambda+cE_{\max}$:
  • Case 1: If $\mathbb{E}_{Y^{\prime}}[f(y+Y^{\prime})]\geq\lambda+cE_{\max}$, then for all $z\geq 0$, the derivative in (34) is non-negative. This implies that the cost function is increasing with respect to $z$, and thus the optimal waiting time is $z^{\star}(y,\mathbf{l};r,\lambda,c)=0$.
  • Case 2: If $\mathbb{E}_{Y^{\prime}}[f(y+Y^{\prime})]<\lambda+cE_{\max}$, the derivative starts negative at $z=0$. Due to the monotonicity of $f(\cdot)$, there exists a unique target value, which we denote momentarily as $t^{\star}$, such that $\mathbb{E}_{Y^{\prime}}[f(t^{\star}+Y^{\prime})]=\lambda+cE_{\max}$. In this regime, $Q(y,\mathbf{l},z,r;\lambda,c)$ decreases until $y+z$ reaches $t^{\star}$, after which it increases. Consequently, the optimal waiting time is chosen such that the sampling occurs when the current age reaches this target, i.e., $z^{\star}=t^{\star}-y$.
We can unify these two cases by defining the route-specific target age $\beta^{\star}(r)$ using the following formulation:
$\beta^{\star}(r)=\inf\left\{t\geq 0:\mathbb{E}_{Y^{\prime}\sim Q_{r}}[f(t+Y^{\prime})]\geq\lambda+cE_{\max}\right\}.$ (35)
This definition covers both scenarios: in Case 2, $\beta^{\star}(r)$ is exactly the unique root $t^{\star}$; in Case 1, the condition holds for $t=0$ (or implies $\beta^{\star}(r)\leq y$), resulting in a target that has already been exceeded. Thus, the optimal policy is described by the target-threshold structure:
$z^{\star}(y,\mathbf{l};r,\lambda,c)=(\beta^{\star}(r)-y)^{+}.$ (36)
@@
$Q(y,\mathbf{l},r)=$ $\;cC_{s}-(\lambda^{\star}+cE_{\max})\left((\beta^{\star}(r)-y)^{+}+\mu_{r}\right)$ (37)
$+\mathbb{E}_{Y^{\prime}\sim Q_{r}}\left[F(y+(\beta^{\star}(r)-y)^{+}+Y^{\prime})-F(y)\right]$
$+\mathbb{E}_{Y^{\prime}\sim Q_{r},\mathbf{l}^{\prime}\sim\boldsymbol{p}}\big[V^{*}(Y^{\prime},\mathbf{l}^{\prime};\lambda^{\star},c)\big],$
where $\mu_{r}=\mathbb{E}[Y^{\prime}]$ is the expected delay of route $r$, and the term $F(y+z^{\star}+Y^{\prime})-F(y)$ captures the exact non-linear age penalty accumulated during the $i$-th interval. With the notation $Q(y,\mathbf{l},r)$, the ACOE turns to:
@@
To analyze the threshold structure of $r^{\star}(y,\mathbf{l})$, the following lemmas discuss some important properties of $Q(y,\mathbf{l},r)$ and $V^{*}(y,\mathbf{l};\lambda^{\star},c)$. We begin with properties that hold for general monotonic non-decreasing penalty functions.
@@
The action-value function $Q(y,\mathbf{l},r)$ is independent of the state $\mathbf{l}$, i.e., $Q(y,\mathbf{l},r)=Q(y,r),\;\forall r\in\mathcal{R}$. Hence, the optimality equation simplifies to:
$V^{*}(y,\mathbf{l};\lambda^{\star},c)=\min_{r\in\mathcal{R}(\mathbf{l})}\{Q(y,r)\}.$ (40)
@@
Even though $V^{*}(y,\mathbf{l};\lambda^{\star},c)$ depends on $\mathbf{l}$, once route $r$ has been selected, the term $\mathbb{E}_{Y^{\prime}\sim Q_{r},\mathbf{l}^{\prime}\sim\boldsymbol{p}}\big[V^{*}(Y^{\prime},\mathbf{l}^{\prime};\lambda^{\star},c)\big]$ in the right-hand side of (37) is independent of $\mathbf{l}$. ∎
@@
For any monotonic non-decreasing penalty function $f(\cdot)$, the following assertions hold true:
@@
We now refine our analysis for the case of ordered expected penalties, as defined in (15). In this scenario, the ordering of the action-value derivatives is strictly determined by the route mean delays.
@@
Under the expected penalty ordering condition (15), for any two routes $j,k\in\mathcal{N}$ such that $\mu_{j}>\mu_{k}$, we have
$\frac{\partial Q(y,j)}{\partial y}\geq\frac{\partial Q(y,k)}{\partial y},\quad\forall y\in\mathbb{R}^{+}.$ (41)
@@
See Appendix C. ∎
With Lemma 6 established, we can derive the threshold-based structure of the optimal routing policy for these specific penalty functions.
@@
Consider $N$ routes with their mean delays satisfying $\mu_{1}\geq\mu_{2}\cdots\geq\mu_{N}$. Under the expected penalty ordering condition (15), if route $j$ is optimal at $y=y^{\star}$ for a given availability $\mathbf{l}$, then:
$r^{\star}(y,\mathbf{l})=\operatorname*{arg\,min}_{r\in\mathcal{R}(\mathbf{l})}\{Q(y,r)\}\geq j,\quad\text{if }y>y^{\star},$ (42)
$r^{\star}(y,\mathbf{l})=\operatorname*{arg\,min}_{r\in\mathcal{R}(\mathbf{l})}\{Q(y,r)\}\leq j,\quad\text{if }y<y^{\star}.$
@@
Lemma 7 implies that under the Expected Penalty Ordering property, $r^{\star}(y,\mathbf{l})$ is a monotonically non-decreasing step function with respect to $y$. Consequently, the optimal sampling policy $z^{\star}(y)=(\beta^{\star}(r^{\star}(y,\mathbf{l}))-y)^{+}$ also follows a piecewise structure dictated by the changes in the optimal route.
Finally, we prove that there are at most $\binom{N}{2}$ unique routing thresholds for these specific cases. For any pair of routes $a,b\in\mathcal{N}$ with $a<b$, let $\mathcal{T}_{a,b}$ be the set of ages $y$ satisfying $Q(y,a)=Q(y,b)$. By Lemma 6 and Lemma 7, if such a crossing exists, it is unique (i.e., $|\mathcal{T}_{a,b}|\leq 1$). Thus, we simplify the notation by defining the scalar $\tau_{a,b}$ as the unique element of this set. Since there are at most $\binom{N}{2}=\frac{N(N-1)}{2}$ such pairs, the optimal routing policy is characterized by at most $\binom{N}{2}$ switching thresholds. Note that the existence of a threshold $\tau_{a,b}$ does not necessitate its presence in the optimal solution.
For general monotonically non-decreasing penalty functions allowing at most $m$ crossover points between any pair of routes, the policy is characterized by at most $m\binom{N}{2}$ thresholds. We expect neither $m$ nor the overall number of thresholds to be large for practical penalty functions and delay distributions.
VII Simulation Results